you have $12000 to invest.if part is invested at 10% and rest at 15%,how should be invested ateach rate to yeild 12% on total amount?
Answer by checkley77(12844)

.15x+.10(12,000-x)=.12*12,000
.15x+1,200-.10x=1440
.05x=1440-1200
.05x=240
x=240/.05
x=4,800 invested @ 15%.
12,000-4,800=7,200 invested @ 10%.Proof:
.15*4,800+.10*7,200=.12*12,000
720+720=1,440
1,440=1,440